Consider the country of Morales with the following closed economy conditions:

 

Y = C + I + G     where:

 

C = 100 + 0.75Y

I = 180 – 15i

G = 100

T = 80

Md = 0.2Y – 5i

Ms = 85


  1. Determine the IS and LM relations 
  2. Find out the equilibrium income Y and Interest Rate i 
  3. Suppose the government increases its purchase by 60 (ΔG = 60)(4 points – show all mathematical work). You will need to redo all the process with the new purchase value which the government has done
  4. Economic analysis: Make a comparative analysis of both national outputs and interest rates – no words limit 
1
Expert's answer
2021-03-16T08:55:59-0400

1. (a)IS relation.

"Y=C+I+G"

"Y= 100+0.75Y+180-15i+100"

"0.75Y=100+180+100-15i"

"Y=506.67-20i" : Which is IS relations.


(b) LM relation.

"\left (M^{\smash {d}}\right)=\left (M^{\smash {s}}\right)"

"0.25Y-5i = 85"

"5i=85-0.25Y"

"i=17-0.05Y": Which is the LM relation.


2.

IS: "Y=506.67-20i"

LM: "i=17-0.05Y"

We first subtitute i in IS to get the value of Y.

"Y=506.67-20 (17-0.05Y)"

"Y=506.67-340"

"Y=166.67"

We now subtitute Y in LM to get the value of i.

"i=17-0.05Y"

"i=17-0.05 (166.67)"

"i=8.67%"

(3.) Initially, G was 250. With increase of 60, the new G is 310.

New Y"=100+0.75Y+180-15i+160"

"Y=586.67-20i"

"Y=246.67"

The new income rate will be:

"i=17-0.05 (246.67)"

"i=4.67"

(4). The new equilibrium gives us lower interest rate of 4.67 compared to the initial 8.67%.